# Edward Witten

### v1

@@ 6.5 Negative Cosmological Constant
Since there is no black hole solution for $T<1/\beta_{0}$, how do we describe thermodynamics in Anti de Sitter space at low temperatures? The answer is straightforward: at low temperatures, we consider an ordinary thermal ensemble of particles in AdS space, described by perturbation theory around the classical solution ${\rm AdS}_{\beta}$. For $\beta>\beta_{0}$, this is presumed to be the dominant or perhaps only classical solution that contributes to a thermal ensemble at the given value of $\beta$. For $\beta<\beta_{0}$, there are two classical solutions to consider, namely ${\rm AdS}_{\beta}$ and the black hole with $r_{+}>r_{0}$. We expect that in the limit $G\to 0$, the thermal ensemble is dominated by the solution of lower action. The difference between the actions of the two solutions is precisely what we computed in eqn. (115). The black hole solution dominates if and only if this difference of actions is negative, which is so precisely if $r_{+}>b$. Therefore, for $r_{+}>b$, the black hole is stable; a thermal gas without a black hole, described by the ${\rm AdS}_{\beta}$ solution, is stable locally at this temperature but should be able to tunnel to a state that contains a black hole. [^21] An explicit description of this tunneling is unknown. For $r_{+}<b$, the thermal gas is stable. For $\frac{b}{\sqrt{3}}<r_{+}<b$, the black hole solution has positive action, so the black hole is stable locally, but can potentially tunnel to thermal ${\rm AdS}_{\beta}$. The phase transition between thermal AdS and the black hole at $r_{+}=b$ is called the Hawking-Page transition. Of course, this is only a true sharp phase transition in the limit $G\to 0$.
@@ 6.6 Positive Cosmological Constant
The portion of de Sitter space that this coordinate system describes is precisely the right static patch ${\mathcal{P}}_{r}$. The observer worldline is at $r=0$, and the coordinate $\tau$ is the observer’s proper time. The observer’s horizon is in this description at $r=\varrho$, with any $\tau$, and is once again a two-sphere of radius $\varrho$. (The $r,\tau$ coordinate system actually breaks down at the horizon and does not describe the portion of the spacetime beyond the horizon.) This way to write the metric of de Sitter space hopefully exhibits an analogy between the de Sitter cosmological horizon at $r=\varrho$ and the Schwarzschild cosmological horizon at $r=2GM$. The difference is that in the case of a black hole, we usually consider an observer living “outside” the cosmological horizon, while in the static patch metric (129), the observer lives at $r=0$ and is surrounded by the cosmological horizon. The boost generator $H$ of the $X_{1}-X_{2}$ plane corresponds in this coordinate system to a symmetry that advances the observer’s proper time
@@
Analytic continuation of correlators from Euclidean signature has given a very quick and efficient way to describe the state $\Psi^{\prime}_{\rm dS}$. This characterization of the state is also illuminating, as it makes manifest the invariance of the state under the de Sitter symmetry $SO(4,1)$. However, it is important to also understand how to describe the state as a functional of the quantum fields on a suitable spacelike hypersurface. Apart from other applications that will become clear, this will fill a serious gap in the definition of $\Psi^{\prime}_{\rm dS}$. We described this state by a set of correlation functions, but to show that there really is a quantum state with those correlation functions, we need to verify a positivity condition $\langle\Psi^{\prime}|{\mathcal{O}}^{\dagger}{\mathcal{O}}|\Psi^{\prime}\rangle\geq 0$, where ${\mathcal{O}}$ is a general linear combination of products $\phi(x_{1})\phi(x_{2})\cdots\phi(x_{s})$. This is obvious from an alternative definition of $\Psi^{\prime}_{\rm dS}$ that we explain next.
@@ 6.7 The Thermofield Double State of a Black Hole
If we do precisely this in the case of a black hole, we run into a snag. The problem can be seen if we recall that the near horizon geometry of the Euclidean black hole, after changing variables from $r$ to a suitable variable $\widetilde{r}$ that vanishes on the horizon is (as in eqn. (78)
${\mathrm{d}}s^{2}=\frac{\widetilde{r}^{2}t_{\sf E}^{2}}{(4GM)^{2}}+{\mathrm{d}}\widetilde{r}^{2}.$ (157)
plus the metric of a two-sphere of radius $2GM$. Eqn. (157) describes ${\mathbb{R}}^{2}$ in polar coordinates if $t_{\sf E}$ is regarded as a polar angle of period $8\pi GM$. Thus the condition $t_{\sf E}=0$ defines a ray in ${\mathbb{R}}^{2}$, ending at the horizon, which is the origin of polar coordinates at $t_{\sf E}=0$. Hence an attempt to define a Cauchy hypersurface by a condition $t_{\sf E}=0$ will fail: this condition defines a partial Cauchy hypersurface, a manifold with a boundary at the horizon.
Instead, we should slightly modify the approach and define ${\mathcal{S}}$ to be the fixed point set of the time-reversal symmetry $t\leftrightarrow-t$. As we noted in section 6.1, this symmetry is the reason that the Schwarzschild solution has a continuation to Euclidean signature. In Lorentz signature, the fixed point set of $t\leftrightarrow-t$ is simply the hypersurface $t=0$. In Euclidean signature, the symmetry $t\leftrightarrow-t$ becomes $t_{\sf E}\leftrightarrow-t_{\sf E}$. Since $t_{\sf E}$ is a periodic variable of period $\beta_{\rm H}=8\pi GM$, the transformation $t_{\sf E}\leftrightarrow-t_{\sf E}$ actually has fixed points both at $t_{\sf E}=0$ and at $t_{\sf E}=\frac{1}{2}\beta_{\rm H}$. In the copy of ${\mathbb{R}}^{2}$ described in eqn. (157), the two rays $t_{\sf E}=0$ and $t_{\sf E}=\frac{1}{2}\beta_{\rm H}$ fit together into a copy of ${\mathbb{R}}$ that actually is a geodesic. This is no coincidence; the fixed point set of a symmetry of a manifold is always a submanifold, never a manifold with boundary. The problem with defining a submanifold by $t_{\sf E}=0$ was that $t_{\sf E}$ fails to be a good coordinate at the horizon.
@@
We decompose the initial value surface ${\mathcal{S}}$ as ${\mathcal{S}}_{\ell}\cup{\mathcal{S}}_{r}$, where ${\mathcal{S}}_{r}$ is the portion at $U<0,\,V>0$, and ${\mathcal{S}}_{\ell}$ is the portion at $U>0,\,V<0$. The domain of dependence of ${\mathcal{S}}_{r}$ is the whole outside-the-horizon wedge ${\mathcal{U}}_{r}$, and the domain of dependence of ${\mathcal{S}}_{\ell}$ is the whole outside-the-horizon wedge ${\mathcal{U}}_{\ell}$. Associated to this, we formally split [^33] The usual remark of footnote 15 applies again. The splitting claimed in the text are not strictly valid in continuum quantum field theory. Tomita-Takesaki theory supplies a rigorous approach. (In the context of quantum gravity, as opposed to quantum field theory in a fixed background spacetime, the splitting $H=H_{r}-H_{\ell}$ is rigorous if $H_{r}$ and $H_{\ell}$ are interpreted as the ADM masses at infinity.) the Hilbert space as
@@ 7.1 Von Neumann Entropy
It can be very hard to know if a system that appears to be in thermal equilibrium is truly in a thermal state with $S_{\rm vN}=S_{\rm th}$, or is really in a state in which $S_{\rm vN}\ll S_{\rm th}$ with only one or a relatively small number of the $p_{i}$ being nonzero. That is because, since the states $|i\rangle$ in this discussion are by hypothesis microstates consistent with the macroscopic properties of the system, they are hard to distinguish. A strong version of this statement is the Entanglement Thermalization Hypothesis (ETH), which says that for a generic macroscopic quantum system, generic microstates (usually taken to be energy eigenstates) cannot be distinguished from a thermal ensemble by any reasonably simple measurement ETH ; ETH2 . A simple measurement is, for example, a measurement of an $n$-point correlation function, where $n$ is a fixed number independent of the number of atoms in the system under study (which, near the thermodynamic limit, is presumed to be much larger than $n$).
@@ 7.3 The Page Curve
$S_{\rm vN}({\sf C})=S_{\rm vN}({\rm BH})$ when the black hole is born, as ${\sf C}$ and the star were in a pure state, and the collapse of the star to a black hole was unitary.
@@ 8.1 Bekenstein-Hawking Entropy as Entanglement Entropy
where $f$ is a constant and $\Lambda$ is an ultraviolet cutoff. On the other hand, in Hawking’s original calculation of black hole evaporation (and in our sketch of this calculation in section 3), loop effects that renormalize Newton’s constant were not taken into account. This means that we should think of the “$G$” that appears in that analysis as a bare Newton constant )$G_{0}$, and thus we should have written the generalized entropy as
@@ 8.2 A Sample Computation
Of course, one question here is whether the analytic continuation away from integer values of $n$ exists and is unique. In ordinary quantum mechanics, one can address this question as follows. As $\rho$ is a positive operator with trace 1, it follows that in the half-plane ${\rm Re}\,n\geq 1$, the function $f(n)={\rm Tr}\,\rho^{n}$ is holomorphic and bounded by $|f(n)|\leq 1$. By Carlson’s theorem (see for example Boas , p. 153), a holomorphic function in that half-plane that coincides with $f(n)$ at positive integers is unique, under condition of being bounded or even under a much weaker condition that allows some exponential growth. So in ordinary quantum mechanics, there is no difficulty with the existence and uniqueness of the analytic continuation away from positive integer values of $n$. In quantum field theory, these considerations are not really applicable, as the formalism of density matrices is not rigorous in quantum field theory. [^40] The factorization of Hilbert space that is usually used in defining density matrices is not valid in continuum quantum field theory, as previously remarked in footnote 15. See also a remark at the end of this section. But in practice, the use of the replica trick to compute entropies in quantum field theory has been very effective.
@@
A slightly more abstract description is as follows. If $P$ be the operator that cyclically permutes the $n$ replicas, then
@@
The upshot of all this is that ${\rm Tr}\,\rho^{n}$ is given by the path integral of the CFT under study on $C_{n}$. However, there is a simpler approach. Away from the branch points, $C_{n}$ is just $n$ copies of the original complex $z$-plane $C$. Instead of studying one copy of theory ${\mathcal{T}}$ on an $n$-fold cover of $C$, it is equivalent, away from the branch points, to study $n$ copies of theory ${\mathcal{T}}$ on $C$. We will denote as ${\mathcal{T}}^{n}$ the CFT that consists of $n$ copies of theory ${\mathcal{T}}$. We would like to compute ${\rm Tr}\,\rho^{n}$ by studying theory ${\mathcal{T}}^{n}$ on $C$. In doing this, what are we supposed to say about the branch points of the covering map $C_{n}\to C$? This question actually has a simple answer. Theory ${\mathcal{T}}^{n}$ is invariant under the group of permutations of the $n$ copies. Under broad conditions, [^41] The discrete symmetry must be one that could be gauged; its ’t Hooft anomaly must vanish. That condition is satisfied in the present example. to a discrete symmetry $\gamma$ of a CFT, one can associate a “twist field” $\Theta_{\gamma}$, with the property that in going around a point $p$ at which the operator $\Theta_{\gamma}(p)$ is inserted, the fields undergo the automorphism $\gamma$. Such twist fields were first constructed in DFMS and have relatively simple properties; in particular, the twist field of lowest possible dimension for a given $\gamma$ is a conformal primary. In our problem, at $z=0$, we want a lowest energy conformal primary [^42] In a lattice regularization, as introduced shortly, one would meet here a non-universal linear combination of all possible twist fields for given $\gamma$. But in the limit that the lattice scale $\varepsilon$ becomes small, the dominant contribution will come from the twist field for $\gamma$ of lowest possible dimension. That is why the relevant twist field $\Theta_{(n)}$ is the one of lowest dimension, which in particular is a conformal primary. There are many other conformal primary twist fields of higher energy for the same $\gamma$. twist field $\Theta_{(n)}$ that cyclically permutes the $n$ sheets; at $z=L$, we need the conjugate twist field $\overline{\Theta}_{(n)}$ that cyclically permutes the $n$ sheets in the opposite direction.
@@ 8.3 The Bekenstein Bound
Let us first discuss the first term ${\rm Tr}\,\rho\log\rho-{\rm Tr}\,\sigma\log\sigma$. Formally, this is a difference of entropies: $-{\rm Tr}\,\sigma\log\sigma$ is $S_{\rm vN}(\sigma)$, the von Neumann entropy of the density matrix $\sigma$, and similarly ${\rm Tr}\,\rho\log\rho$ is $-S_{\rm vN}(\rho)$, the negative of the von Neumann entropy of $\rho$. In a Hamiltonian approach, these von Neumann entropies measure entanglement in the states $\Psi$ or $\Omega$ between modes in the partial Cauchy hypersurface ${\mathcal{S}}_{r}$ and modes in the complementary partial Cauchy hypersurface ${\mathcal{S}}_{\ell}$; in a covariant description, they measures entanglement between modes in the Rindler wedge $\mathcal{R}_{r}$ and modes in the complementary Rindler wedge $\mathcal{R}_{\ell}$. The von Neumann entropy is this problem is ultraviolet divergent because of entanglement between short wavelength modes close to but on on opposite sides of the common boundary of ${\mathcal{S}}_{\ell}$ and ${\mathcal{S}}_{r}$ at $x=0$. That common boundary, which is often called the entangling surface, is is also the “edge” or corner at which $R_{\ell}$ and $R_{r}$ meet; it was denoted as $\Sigma$ in fig. 7. We illustrated the ultraviolet divergence in the entanglement entropy in a concrete example in section 8.2. Because every state looks like the vacuum at short distances, the ultraviolet divergence is independent of the state and therefore the difference $\Delta S=S_{\rm vN}(\rho)-S_{\rm vN}(\sigma)$ is ultraviolet finite. Now we can understand one contribution on the right hand side of eqn. (241): it is precisely $-\Delta S$.

### v4

@@ 6.5 Negative Cosmological Constant
Since there is no black hole solution for $T<1/\beta_{0}$, how do we describe thermodynamics in Anti de Sitter space at low temperatures? The answer is straightforward: at low temperatures, we consider an ordinary thermal ensemble of particles in AdS space, described by perturbation theory around the classical solution ${\rm AdS}_{\beta}$. For $\beta>\beta_{0}$, this is presumed to be the dominant or perhaps only classical solution that contributes to a thermal ensemble at the given value of $\beta$. For $\beta<\beta_{0}$, there are two classical solutions to consider, namely ${\rm AdS}_{\beta}$ and the black hole with $r_{+}>r_{0}$. We expect that in the limit $G\to 0$, the thermal ensemble is dominated by the solution of lower action. The difference between the actions of the two solutions is precisely what we computed in eqn. (115). The black hole solution dominates if and only if this difference of actions is negative, which is so precisely if $r_{+}>b$. Therefore, for $r_{+}>b$, the black hole is stable; a thermal gas without a black hole, described by the ${\rm AdS}_{\beta}$ solution, is stable locally at this temperature but should be able to tunnel to a state that contains a black hole. [^21] An explicit description of this tunneling is unknown. For $r_{+}<b$, the thermal gas is stable. For $\frac{b}{\sqrt{3}}<r_{+}<b$, the action of the black hole solution is bigger, so although the black hole is stable locally, it can potentially tunnel to thermal ${\rm AdS}_{\beta}$. The phase transition between thermal AdS and the black hole at $r_{+}=b$ is called the Hawking-Page transition. Of course, this is only a true sharp phase transition in the limit $G\to 0$.
@@ 6.6 Positive Cosmological Constant
The portion of de Sitter space that this coordinate system describes is precisely the right static patch ${\mathcal{P}}_{r}$. The observer worldline is at $r=0$, and the coordinate $\tau$ is the observer’s proper time. The observer’s horizon is in this description at $r=\varrho$, with any $\tau$, and is once again a two-sphere of radius $\varrho$. (The $r,\tau$ coordinate system actually breaks down at the horizon and does not describe the portion of the spacetime beyond the horizon.) This way to write the metric of de Sitter space hopefully exhibits an analogy between the de Sitter cosmological horizon at $r=\varrho$ and the Schwarzschild cosmological horizon at $r=2GM$. The difference is that in the case of a black hole, we usually consider an observer living “outside” the horizon, while in the static patch metric (129), the observer lives at $r=0$ and is surrounded by the cosmological horizon. The boost generator $H$ of the $X_{1}-X_{2}$ plane corresponds in this coordinate system to a symmetry that advances the observer’s proper time
@@
Analytic continuation of correlators from Euclidean signature has given a very quick and efficient way to describe the state $\Psi^{\prime}_{\rm dS}$. This characterization of the state is also illuminating, as it makes manifest the invariance of the state under the de Sitter symmetry $SO(4,1)$. However, it is important to also understand how to describe the state as a functional of the quantum fields on a suitable spacelike hypersurface. Apart from other applications that will become clear, this will fill a serious gap in the definition of $\Psi^{\prime}_{\rm dS}$. We described this state by a set of correlation functions, but to show that there really is a quantum state with those correlation functions, we need to verify a positivity condition $\langle\Psi^{\prime}_{\rm dS}|{\mathcal{O}}^{\dagger}{\mathcal{O}}|\Psi^{\prime}_{\rm dS}\rangle\geq 0$, where ${\mathcal{O}}$ is a general linear combination of products $\phi(x_{1})\phi(x_{2})\cdots\phi(x_{s})$. This is obvious from an alternative definition of $\Psi^{\prime}_{\rm dS}$ that we explain next.
@@ 6.7 The Thermofield Double State of a Black Hole
If we do precisely this in the case of a black hole, we run into a snag. The problem can be seen if we recall that the near horizon geometry of the Euclidean black hole, after changing variables from $r$ to a suitable variable $\widetilde{r}$ that vanishes on the horizon is (as in eqn. (78))
${\mathrm{d}}s^{2}=\frac{\widetilde{r}^{2}{\mathrm{d}}t_{\sf E}^{2}}{(4GM)^{2}}+{\mathrm{d}}\widetilde{r}^{2}.$ (157)
plus the metric of a two-sphere of radius $2GM$. Eqn. (157) describes ${\mathbb{R}}^{2}$ in polar coordinates if $t_{\sf E}$ is regarded as a polar angle of period $8\pi GM$. Thus the condition $t_{\sf E}=0$ defines a ray in ${\mathbb{R}}^{2}$, ending at the horizon, which is the origin of polar coordinates at $\widetilde{r}=0$. Hence an attempt to define a Cauchy hypersurface by a condition $t_{\sf E}=0$ will fail: this condition defines a partial Cauchy hypersurface, a manifold with a boundary at the horizon.
Instead, we should slightly modify the approach and define ${\mathcal{S}}$ to be the fixed point set of the time-reversal symmetry $t\leftrightarrow-t$. As we noted in section 6.1, this symmetry is the reason that the Schwarzschild solution has a continuation to Euclidean signature. In Lorentz signature, the fixed point set of $t\leftrightarrow-t$ is simply the hypersurface $t=0$. In Euclidean signature, the symmetry $t\leftrightarrow-t$ becomes $t_{\sf E}\leftrightarrow-t_{\sf E}$. Since $t_{\sf E}$ is a periodic variable of period $\beta_{\rm H}=8\pi GM$, the transformation $t_{\sf E}\leftrightarrow-t_{\sf E}$ actually has fixed points both at $t_{\sf E}=0$ and at $t_{\sf E}=\frac{1}{2}\beta_{\rm H}$. In the copy of ${\mathbb{R}}^{2}$ described in eqn. (157), the two rays $t_{\sf E}=0$ and $t_{\sf E}=\frac{1}{2}\beta_{\rm H}$ fit together into a copy of ${\mathbb{R}}$ that actually is a geodesic. This is no coincidence; the fixed point set of a symmetry of a manifold is always a submanifold, never a submanifold with boundary. The problem with defining a submanifold by $t_{\sf E}=0$ was that $t_{\sf E}$ fails to be a good coordinate at the horizon.
@@
We decompose the initial value surface ${\mathcal{S}}$ as ${\mathcal{S}}_{\ell}\cup{\mathcal{S}}_{r}$, where ${\mathcal{S}}_{r}$ is the portion at $U<0,\,V>0$, and ${\mathcal{S}}_{\ell}$ is the portion at $U>0,\,V<0$. The domain of dependence of ${\mathcal{S}}_{r}$ is the whole outside-the-horizon wedge ${\mathcal{U}}_{r}$, and the domain of dependence of ${\mathcal{S}}_{\ell}$ is the whole outside-the-horizon wedge ${\mathcal{U}}_{\ell}$. Associated to this, we formally split [^33] The usual remark of footnote 15 applies again. The splitting claimed in the text are not strictly valid in continuum quantum field theory. Tomita-Takesaki theory supplies a rigorous approach. (In the context of quantum gravity, as opposed to quantum field theory in a fixed background spacetime, the splitting $H=H_{r}-H_{\ell}$ is rigorous if $H_{r}$ and $H_{\ell}$ are interpreted as the ADM energies at infinity.) the Hilbert space as
@@ 7.1 Von Neumann Entropy
It can be very hard to know if a system that appears to be in thermal equilibrium is truly in a thermal state with $S_{\rm vN}=S_{\rm th}$, or is really in a state in which $S_{\rm vN}\ll S_{\rm th}$ with only one or a relatively small number of the $p_{i}$ being nonzero. That is because, since the states $|i\rangle$ in this discussion are by hypothesis microstates consistent with the macroscopic properties of the system, they are hard to distinguish. A strong version of this statement is the Eigenvalue Thermalization Hypothesis (ETH), which says that for a generic macroscopic quantum system, generic microstates (usually taken to be energy eigenstates) cannot be distinguished from a thermal ensemble by any reasonably simple measurement ETH ; ETH2 . A simple measurement is, for example, a measurement of an $n$-point correlation function, where $n$ is a fixed number independent of the number of atoms in the system under study (which, near the thermodynamic limit, is presumed to be much larger than $n$).
@@ 7.3 The Page Curve
Moreover, $S_{\rm vN}({\sf C})=S_{\rm vN}({\rm BH})$ when the black hole is born, as ${\sf C}$ and the star were in a pure state, and the collapse of the star to a black hole was unitary.
@@ 8.1 Bekenstein-Hawking Entropy as Entanglement Entropy
where $f$ is a constant and $\Lambda$ is an ultraviolet cutoff. On the other hand, in Hawking’s original calculation of black hole evaporation (and in our sketch of this calculation in section 3), loop effects that renormalize Newton’s constant were not taken into account. This means that we should think of the “$G$” that appears in that analysis as a bare Newton constant $G_{0}$, and thus we should have written the generalized entropy as
@@ 8.2 A Sample Computation
Of course, one question here is whether the analytic continuation away from integer values of $n$ exists and is unique. In ordinary quantum mechanics, one can address this question as follows. As $\rho$ is a positive operator with trace 1, it follows that in the half-plane ${\rm Re}\,n\geq 1$, the function $f(n)={\rm Tr}\,\rho^{n}$ is holomorphic and bounded by $|f(n)|\leq 1$. So the analytic continuation exists. The continuation is unique by Carlson’s theorem. According to this theorem (see for example Boas , p. 153), a holomorphic function in the half-plane ${\rm Re}\,n\geq 1$ that coincides with $f(n)$ at positive integers is unique, under condition of being bounded or even under a much weaker condition that allows some exponential growth. So in ordinary quantum mechanics, there is no difficulty with the existence and uniqueness of the analytic continuation away from positive integer values of $n$. In quantum field theory, these considerations are not really applicable, as the formalism of density matrices is not rigorous in quantum field theory. [^40] The factorization of Hilbert space that is usually used in defining density matrices is not valid in continuum quantum field theory, as previously remarked in footnote 15. See also a remark at the end of this section. But in practice, the use of the replica trick to compute entropies in quantum field theory has been very effective.
@@
A slightly more abstract description is as follows. If $P$ is the operator that cyclically permutes the $n$ replicas, then
@@
The upshot of all this is that ${\rm Tr}\,\rho^{n}$ is given by the path integral of the CFT under study on $C_{n}$. However, there is a simpler approach. Away from the branch points, $C_{n}$ is just $n$ copies of the original complex $z$-plane $C$. Instead of studying one copy of theory ${\mathcal{T}}$ on an $n$-fold cover of $C$, it is equivalent, away from the branch points, to study $n$ copies of theory ${\mathcal{T}}$ on $C$. We will denote as ${\mathcal{T}}^{n}$ the CFT that consists of $n$ copies of theory ${\mathcal{T}}$. We would like to compute ${\rm Tr}\,\rho^{n}$ by studying theory ${\mathcal{T}}^{n}$ on $C$. In doing this, what are we supposed to say about the branch points of the covering map $C_{n}\to C$? This question actually has a simple answer. Theory ${\mathcal{T}}^{n}$ is invariant under the group of permutations of the $n$ copies. Under broad conditions, [^41] The discrete symmetry must be one that could be gauged; its ’t Hooft anomaly must vanish. That condition is satisfied in the present example. to a discrete symmetry $\gamma$ of a CFT, one can associate a “twist field” $\Theta_{\gamma}$, with the property that in going around a point $p$ at which the operator $\Theta_{\gamma}(p)$ is inserted, the fields undergo the automorphism $\gamma$. Such twist fields were first constructed in DFMS and have relatively simple properties; in particular, the twist field of lowest possible dimension for a given $\gamma$ is a conformal primary. In our problem, at $z=0$, we want a lowest energy conformal primary [^42] In a lattice regularization, as introduced shortly, one would meet here a non-universal linear combination of all possible twist fields for given $\gamma$. But in the limit that the lattice scale $\varepsilon$ becomes small, the dominant contribution will come from the twist field for $\gamma$ of lowest possible dimension. That is why the relevant twist field $\Theta_{(n)}$ is the one of lowest dimension, which in particular is a conformal primary. There are many other conformal primary twist fields of higher dimension for the same $\gamma$. twist field $\Theta_{(n)}$ that cyclically permutes the $n$ sheets; at $z=L$, we need the conjugate twist field $\overline{\Theta}_{(n)}$ that cyclically permutes the $n$ sheets in the opposite direction.
@@ 8.3 The Bekenstein Bound
Let us first discuss the first term ${\rm Tr}\,\rho\log\rho-{\rm Tr}\,\sigma\log\sigma$. Formally, this is a difference of entropies: $-{\rm Tr}\,\sigma\log\sigma$ is $S_{\rm vN}(\sigma)$, the von Neumann entropy of the density matrix $\sigma$, and similarly ${\rm Tr}\,\rho\log\rho$ is $-S_{\rm vN}(\rho)$, the negative of the von Neumann entropy of $\rho$. In a Hamiltonian approach, these von Neumann entropies measure entanglement in the states $\Psi$ or $\Omega$ between modes in the partial Cauchy hypersurface ${\mathcal{S}}_{r}$ and modes in the complementary partial Cauchy hypersurface ${\mathcal{S}}_{\ell}$; in a covariant description, they measures entanglement between modes in the Rindler wedge $\mathcal{R}_{r}$ and modes in the complementary Rindler wedge $\mathcal{R}_{\ell}$. The von Neumann entropy is this problem is ultraviolet divergent because of entanglement between short wavelength modes close to but on opposite sides of the common boundary of ${\mathcal{S}}_{\ell}$ and ${\mathcal{S}}_{r}$ at $x=0$. That common boundary, which is often called the entangling surface, is also the “edge” or corner at which $R_{\ell}$ and $R_{r}$ meet; it was denoted as $\Sigma$ in fig. 7. We illustrated the ultraviolet divergence in the entanglement entropy in a concrete example in section 8.2. Because every state looks like the vacuum at short distances, the ultraviolet divergence is independent of the state and therefore the difference $\Delta S=S_{\rm vN}(\rho)-S_{\rm vN}(\sigma)$ is ultraviolet finite. Now we can understand one contribution on the right hand side of eqn. (241): it is precisely $-\Delta S$.